# LDP for invariant measures of the Allen-Cahn equation with inhomogeneous boundary conditions and multiplicative noise

### v1

@@ 3.1 Energy functional and quasi-potential
In this section, we study the rate functional $U_{L}$ for invariant measures. We first introduce the energy functional $\mathbf{E}_{L}$. For functions $u:[-L,L]\rightarrow\mathbb{R}$ with boundary conditions $u(\pm L)=\pm 1$, we consider the energy functional
$\mathbf{E}_{L}(u):=\int_{-L}^{L}\left[\frac{|u^{\prime}(\xi)|^{2}}{2}+\frac{1}{4}(u^{2}(\xi)-1)^{2}\right]\;d\xi.$ (3.1)
@@
$\begin{split}\mathbf{E}_{L}(u)&=\int_{-L}^{L}\frac{|u^{\prime}(\xi)|^{2}}{2}+\frac{1}{4}(u^{2}(\xi)-1)^{2}\;d\xi\\ &=\int_{-L}^{L}\Bigg[\frac{1}{2}\left(u^{\prime}(\xi)-\sqrt{\frac{1}{2}(u^{2}(\xi)-1)^{2}+e_{L}}\right)^{2}\\ &\ \ \ \ +\sqrt{\frac{1}{2}(u^{2}(\xi)-1)^{2}+e_{L}}\cdot u^{\prime}(\xi)-\frac{1}{2}e_{L}\Bigg]\;d\xi\\ &=\frac{1}{2}\int_{-L}^{L}\left(u^{\prime}(\xi)-\sqrt{\frac{1}{2}(u^{2}(\xi)-1)^{2}+e_{L}}\right)^{2}\;d\xi\\ &\ \ \ \ +\int_{-1}^{1}\sqrt{\frac{1}{2}(u^{2}-1)^{2}+e_{L}}\;du-Le_{L}.\end{split}$ (3.4)
@@
so that $\mathbf{E}_{L}^{*}(m_{L}-\psi)=0$. Note that $m_{L}-\psi$ is the only solution for equation
@@
The following proposition is needed in this paper:
@@
For $\bar{u}\in E$, we denote $u=\bar{u}+\psi$. Assume $\mathbf{E}_{L}^{*}(\bar{u})\leq\eta$, according to (3.3), (3.5) and (3.6),
@@
where $V(\cdot)=\sqrt{\frac{1}{2}((\cdot)^{2}-1)^{2}+e_{L}}$ is a locally Lipchitz, and $\phi$ satisfies $\int_{-L}^{L}\phi^{2}(\xi)\;d\xi\leq 2\eta$. By the Cauchy-Schwarz inequality,
@@
This implies we can choose $\eta$ sufficiently small such that $\sup_{\xi\in[-L,L]}|u(\xi)|\leq 2$. Thus, by (3.3) and the locally Lipchitz property of $V(\cdot)$, one obtains that
@@
By Gronwall’s inequality, one can choose $\eta$ small enough such that
$\|\bar{u}-(m_{L}-\psi)\|_{E}=\sup_{\xi\in[-L,L]}|u(\xi)-m_{L}(\xi)|\leq 2\sqrt{L\eta}e^{2LC}\leq\varepsilon.$
@@
Note that the level sets of $\mathbf{E}_{L}^{*}$ are compact (see Proposition 3.1 of [bertini2008dobrushin]). By Sobolev embedding theorem, $\mathbf{E}_{L}^{*}(\bar{u})$ is finite if and only if $\bar{u}\in H^{1}(-L,L)$. For $\bar{u},h\in H^{1}(-L,L)$, we denote $D_{h}\mathbf{E}_{L}^{*}(\bar{u})$ to be the Fréchet derivative of $\mathbf{E}_{L}^{*}$ on $\bar{u}$,
@@
Here, the last equality is due to integration by parts and $\Delta\psi=0$. This implies that
@@
The functions $U_{L}$ and $\mathbf{E}_{L}^{*}$ has the following connection.
@@
Suppose $U_{L}(\zeta)=+\infty$, then our claim is clearly true. Therefore, we assume $U_{L}(\zeta)<\infty$. Then, there exists some $T>0$ and $z\in C([0,T];E)$ satisfying $z(0)=m_{L}-\psi$, $z(T)=\zeta$ such that $z$ is a solution of equation (2.3) with $f\in L^{2}(0,T;H)$ such that
@@
Following the same procedure of the proof in Proposition 6.1 of [cerrai2003stochastic], we have
$\sup_{t\geq 0}\|z(t)\|_{E}\leq C\|f\|_{L^{2}([0,+\infty);H)}.$ (3.12)
@@
This implies that there exists $t^{\prime}>0$ such that $z_{0}^{x}(t^{\prime})\in H^{1}(-L,L)$. Multiplying both parts of equation (2.3) by $D\mathbf{E}^{*}_{L}(z^{x}_{0}(t))$ and integrating with respect to $\xi$ in [L,L][-L,L]. Since $f=0$, by chains rule
@@
$\mathbf{E}_{L}^{*}(z_{0}^{x}(t^{\prime}))\geq\mathbf{E}^{*}_{L}(z^{x}_{0}(t))\geq\frac{1}{2}\|z_{0}^{x}(t)+\psi\|_{H^{1}}^{2}-\mathbf{E}_{L}(m_{L})\geq\|z_{0}^{x}(t)\|_{H^{1}}^{2}+\frac{1}{L}-\mathbf{E}_{L}(m_{L}).$
@@
$\lim_{k\rightarrow+\infty}\langle F(z_{0}^{x}(t_{k}))-F(z_{L}),\tilde{h}\rangle_{H}\leq C\lim_{k\rightarrow+\infty}\langle|z_{0}^{x}(t_{k})-z_{L}|,\tilde{h}\rangle_{H}=0.$
@@
$\lim_{k\rightarrow+\infty}\langle\Delta z_{0}^{x}(t_{k})+F(z_{0}^{x}(t_{k})),h\rangle_{H}$ $=\lim_{k\rightarrow+\infty}\left(\langle D_{\xi}z_{0}^{x}(t_{k}),D_{\xi}h\rangle_{H}+\langle F(z_{0}^{x}(t_{k})),h\rangle_{H}\right)$
$=\langle D_{\xi}z_{L},D_{\xi}h\rangle_{H}+\langle F(z_{L}),h\rangle_{H}$
@@
Due to the uniqueness of the solution for equation (3.7), one gets
@@
Here, the Cauchy-Schwarz inequality and the local Lipschitz continuity of $F$ are used repeatedly when the above two limits are taken. Therefore, $z_{0}^{x}(t_{k})$ converges to $m_{L}-\psi$ in $H^{1}(-L,L)$ and $\bigcup_{t\geq 1}\{z_{0}^{x}(t,\cdot)\}$ is precompact in $H^{1}(-L,L)$.
@@
Thus, by Sobolev embedding,
@@
According to Lemma 3.3, (3.16) and Sobolev embedding, for any $1\geq\varepsilon>0$ and $\zeta\in E$, there exists $t^{*}>0$ such that
@@
By the definition, $v$ is a path in $C([0,1];E)$ satisfies $v(0)=m_{L}-\psi$ and $v(1)=z_{0}^{\zeta}(t^{*})$.
Note, under our assumptions, the following more explicit expression for $I_{T}^{x}$ can read as:
@@ 3.2 Compactness of level sets of the quasi-potential
In this section, we will show the compactness of level sets of quasi-potential. For any $t_{1}<t_{2}$ and $z\in C([t_{1},t_{2}];E)$ we define
@@
Similarly, for $z\in C((-\infty,0];E)$ we define
@@
According to our assumption, there exists a sequence $t_{n}\nearrow+\infty$, $C>0$ such that $\|z(-t_{n})\|_{E}\leq C$, $n\in\mathbb{N}$ and
@@
where $z$ is the solution of the equation (2.3). Due to (3.12),
@@
Note that $\|f\|_{L^{2}([-t_{n},0];H)}^{2}=2I_{t_{n}}(z_{n})\leq 2r$, by Gronwall’s inequality, we have for $t\in[-t_{n},0]$
@@
Thus, for any $n\in\mathbb{N}$ there exists $t^{\prime}\in[-t_{n},-t_{n}+1]$ such that $\|z(t^{\prime})\|_{H^{1}}^{2}\leq C(\|z(-t_{n})\|_{H}^{2}+2r+1)$. From (3.1) and Sobolev embedding, $\mathbf{E}^{*}_{L}(z(t^{\prime}))\leq C_{r}$. Then, by (3.11), for any $t\geq t^{\prime}$
@@
Since level sets of $\mathbf{E}^{*}_{L}(\cdot)$ are compact in $E$ ([bertini2008dobrushin]), there exist $x\in E$ and a subsequence still denoted by $z_{i}(0)$ such that
@@
Thanks to Lemma 3.3, for any $\eta>0$, one can choose $t_{0}$ sufficiently large so that $\mathbf{E}^{*}_{L}(z_{0}^{x}(t))\leq\frac{\eta}{4}$. Thus, by Sobolev inequality and (3.24), there exists $N$ such that for any $N\leq i\in\mathbb{N}$,
@@
By (3.22), for any $\eta>0$, there exists $N$ such that for all $i\geq N$, $CI_{t_{0}}(z_{i})\leq\frac{\eta}{2}$. Therefore, one obtains that, for $i$ sufficiently large,
@@
and $C_{r}$ is given in Lemma 3.5. As the level sets of $\mathbf{E}^{*}_{L}(\cdot)$ are compact in $E$, we have $K_{\Lambda,-k,0}$ is compact in C([k,0];E)C([-k,0];E).
@@
We have the following characterization of $U_{L}$. For the sake of simplicity of presentation, we denote
@@
If the infimum is attainable, then we write $I_{-\infty}^{\inf}(x)$ as $I_{-\infty}^{\min}(x)$.
@@
where $z$ is the solution of equation (2.3), starting from the time $t_{n}$. Since $f\in L^{2}(-\infty,0;H)$, for any $\varepsilon>0$, there exists $T_{\varepsilon}\leq 0$ such that $\frac{1}{2}\|f\|_{L^{2}(-\infty,T_{\varepsilon};H)}^{2}\leq\varepsilon$. Denote $z_{f}^{-t_{n},m_{L}-\psi}$ as the solution of equation (2.3) with $z_{f}^{-t_{n},m_{L}-\psi}(-t_{n})=m_{L}-\psi$. Since $m_{L}-\psi$ is the equilibrium of (2.3) with $f=0$, we have $z_{0}^{-t_{n},m_{L}-\psi}(t)=m_{L}-\psi$ for every $t\in[-t_{n},0]$. Then, by (3.24), for $n\in\mathbb{N}$ sufficiently large there exists $T_{\varepsilon}\in[-t_{n},-t_{n}+1]$ such that
@@
The characterization of $U_{L}$ allows us to enhance the compactness of the level sets of $I_{-\infty}(\cdot)$ on $C((-\infty,0];E)$ to the compactness of the level sets of $U_{L}(\cdot)$ on $E$. This is the main result of this section and a key result in the theory of large deviations for invariant measures.
@@
such that $U_{L}(x_{n})=I_{-\infty}(z_{n})$. Since $U_{L}(x_{n})\leq r$, we have that $\{z_{n}\}\subset K_{-\infty}(r)$. By Proposition 3.6, $K_{-\infty}(r)$ is compact. So there exists a subsequence $\{z_{n_{k}}\}\subset\{z_{n}\}$ such that $z_{n_{k}}\rightarrow z\in K_{-\infty}(r)$ in $C((-\infty,0];E)$. In particular, $x_{n_{k}}=z_{n_{k}}(0)\rightarrow z(0)$ in $E$. Thanks to Proposition 3.8, we have $U_{L}(z(0))\leq I_{-\infty}(z)\leq r$. That is to say, $z(0)\in\Phi_{L}(r)$. The proof is complete. ∎
4 LDP for the invariant measures
In this section, we will prove the LDP for $\{\mu_{\epsilon}\}_{\epsilon>0}$ which is the main result of this paper. This is indicated by the LDP of $\{\bar{\mu}_{\epsilon}\}_{\epsilon>0}$, whose relation with $\{\mu_{\epsilon}\}_{\epsilon>0}$ is given by (1.5).

### v2

@@ 3.1 Energy functional and quasi-potential
In this section, we study the rate functional $U_{L}$ for invariant measures. We first introduce the energy functional $\mathbf{E}_{L}$. For functions $u:[-L,L]\rightarrow\mathbb{R}$ with boundary conditions $u(\pm L)=\pm 1$, we consider the energy functional (1.7) with
$V(u)=\frac{1}{4}(u^{2}-1)^{2}.$ (3.1)
@@
$\begin{split}\mathbf{E}_{L}(u)&=\int_{-L}^{L}\left[\frac{|u^{\prime}(\xi)|^{2}}{2}+\frac{1}{4}(u^{2}(\xi)-1)^{2}\right]\;d\xi\\ &=\int_{-L}^{L}\Bigg[\frac{1}{2}\left(u^{\prime}(\xi)-\sqrt{\frac{1}{2}(u^{2}(\xi)-1)^{2}+e_{L}}\right)^{2}\\ &\ \ \ \ +\sqrt{\frac{1}{2}(u^{2}(\xi)-1)^{2}+e_{L}}\cdot u^{\prime}(\xi)-\frac{1}{2}e_{L}\Bigg]\;d\xi\\ &=\frac{1}{2}\int_{-L}^{L}\left(u^{\prime}(\xi)-\sqrt{\frac{1}{2}(u^{2}(\xi)-1)^{2}+e_{L}}\right)^{2}\;d\xi\\ &\ \ \ \ +\int_{-1}^{1}\sqrt{\frac{1}{2}(u^{2}-1)^{2}+e_{L}}\;du-Le_{L}.\end{split}$ (3.4)
@@
so that $\mathbf{E}_{L}^{*}(m_{L}-\psi)=0$. Note that $m_{L}-\psi$ is the only solution for the equation
@@
One can prove the following proposition which will be needed in this paper:
@@
For $\bar{u}\in E$, we denote $u=\bar{u}+\psi$. According to (3.4), (3.5), and (3.6),
@@
where $V(\cdot)=\sqrt{\frac{1}{2}((\cdot)^{2}-1)^{2}+e_{L}}$ is locally Lipschitz. Assume $\mathbf{E}_{L}^{*}(\bar{u})\leq\eta$, then $\phi$ satisfies $\int_{-L}^{L}\phi^{2}(\xi)\;d\xi\leq 2\eta$. By the Cauchy-Schwarz inequality,
@@
This implies that we can choose $\eta$ sufficiently small such that $\sup_{\xi\in[-L,L]}|u(\xi)|\leq 2$. Thus, by (3.3) and the locally Lipchitz property of $V(\cdot)$, one obtains
@@
By Grönwall’s inequality, one can choose $\eta$ small enough such that
$\|\bar{u}-(m_{L}-\psi)\|_{E}=\sup_{\xi\in[-L,L]}|u(\xi)-m_{L}(\xi)|\leq 2\sqrt{L\eta}e^{2LC}.$
@@
Note that the level sets of $\mathbf{E}_{L}^{*}$ are compact (see Proposition 3.1 of [bertini2008dobrushin]). By the Sobolev embedding theorem, $\mathbf{E}_{L}^{*}(\bar{u})$ is finite if and only if $\bar{u}\in H^{1}(-L,L)$. For $\bar{u},h\in H^{1}(-L,L)$, denote by $D_{h}\mathbf{E}_{L}^{*}(\bar{u})$ the Fréchet derivative of $\mathbf{E}_{L}^{*}$ on $\bar{u}$, then
@@
Here, the last equality is due to the integration by parts and the fact that $\Delta\psi=0$. This implies that
@@
The functions $U_{L}$ and $\mathbf{E}_{L}^{*}$ have the following connection.
@@
Suppose $U_{L}(\zeta)=+\infty$, then our claim is clearly true. Therefore, we assume $U_{L}(\zeta)<\infty$. Then, there exist some $T>0$ and $z\in C([0,T];E)$ satisfying $z(0)=m_{L}-\psi$, $z(T)=\zeta$ such that $z$ is a solution of equation (2.3) with $f\in L^{2}(0,T;H)$ such that
@@
Following the same procedure as in the proof of Proposition 6.1 of [cerrai2003stochastic], we have
$\sup_{t\geq 0}\|z(t)\|_{E}\leq C(\|f\|_{L^{2}([0,+\infty);H)}+1).$ (3.12)
@@
This implies that there exists $t^{\prime}>0$ such that $z_{0}^{x}(t^{\prime})\in H^{1}(-L,L)$. Multiplying both parts of equation (2.3) by $D\mathbf{E}^{*}_{L}(z^{x}_{0}(t))$ and integrating with respect to $\xi$ in [L,L][-L,L]. Since $f=0$, by the chain rule
@@
$\mathbf{E}_{L}^{*}(z_{0}^{x}(t^{\prime}))\geq\mathbf{E}^{*}_{L}(z^{x}_{0}(t))\geq\frac{1}{2}\|z_{0}^{x}(t)+\psi\|_{H^{1}}^{2}-\mathbf{E}_{L}(m_{L})\geq\|z_{0}^{x}(t)\|_{H^{1}}^{2}-\frac{1}{L}-\mathbf{E}_{L}(m_{L}).$
@@
$\lim_{k\rightarrow+\infty}\langle F(z_{0}^{x}(t_{k}))-F(z_{L}),\tilde{h}\rangle_{H}\leq C\lim_{k\rightarrow+\infty}\langle|z_{0}^{x}(t_{k})-z_{L}|,|\tilde{h}|\rangle_{H}=0.$
@@
$\lim_{k\rightarrow+\infty}\langle\Delta z_{0}^{x}(t_{k})+F(z_{0}^{x}(t_{k})),h\rangle_{H}$ $=\lim_{k\rightarrow+\infty}\left(\langle\nabla z_{0}^{x}(t_{k}),\nabla h\rangle_{H}+\langle F(z_{0}^{x}(t_{k})),h\rangle_{H}\right)$
$=\langle\nabla z_{L},\nabla h\rangle_{H}+\langle F(z_{L}),h\rangle_{H}$
@@
Due to the uniqueness of the solution for equation (3.7), one obtains
@@
Here, the Cauchy-Schwarz inequality and the local Lipschitz continuity of $F$ are used when the above two limits are taken. Therefore, $z_{0}^{x}(t_{k})$ converges to $m_{L}-\psi$ in $H^{1}(-L,L)$ and $\bigcup_{t\geq 1}\{z_{0}^{x}(t,\cdot)\}$ is precompact in $H^{1}(-L,L)$.
@@
Thus, by the Sobolev embedding,
@@
According to Lemma 3.3, (3.16) and the Sobolev embedding theorem, for any $1\geq\varepsilon>0$ and $\zeta\in E$, there exists a real number $t^{*}>0$ such that
@@
By definition, $v$ is a path in $C([0,1];E)$ that satisfies $v(0)=m_{L}-\psi$ and $v(1)=z_{0}^{\zeta}(t^{*})$.
Note that, under our assumptions, the following more explicit expression for $I_{T}^{x}$ can be read as:
@@ 3.2 Compactness of level sets of the quasi-potential
In this section, we will show the compactness of level sets of quasi-potential. For any $t_{1}<t_{2}$ and $z\in C([t_{1},t_{2}];E)$, define
@@
Similarly, for $z\in C((-\infty,0];E)$, define
@@
According to the assumption of the lemma, there exist a sequence $t_{n}\nearrow+\infty$ and a constant $C>0$ such that $\|z(-t_{n})\|_{E}\leq C$, $n\in\mathbb{N}$ and
@@
where $z$ is the solution of equation (2.3) with $z(-t_{n},\cdot)=z(-t_{n})$. Due to (3.12),
@@
Note that $\|f\|_{L^{2}([-t_{n},0];H)}^{2}=2I_{t_{n}}(z_{n})\leq 2r$, by Grönwall’s inequality, we have for $t\in[-t_{n},0]$
@@
Thus, for any $n\in\mathbb{N}$ there exists $t^{\prime}\in[-t_{n},-t_{n}+1]$ such that $\|z(t^{\prime})\|_{H^{1}}^{2}\leq C(\|z(-t_{n})\|_{H}^{2}+2r+1)$. From (1.7) and the Sobolev embedding, $\mathbf{E}^{*}_{L}(z(t^{\prime}))\leq C_{r}$. Then, by (3.11), for any $t\geq t^{\prime}$
@@
Since the level sets of $\mathbf{E}^{*}_{L}(\cdot)$ are compact in $E$ ([bertini2008dobrushin]), there exist $x\in E$ and a subsequence still denoted by $z_{i}(0)$ such that
@@
Thanks to Lemma 3.3, for any $\eta>0$, one can choose $t_{0}$ sufficiently large so that $\mathbf{E}^{*}_{L}(z_{0}^{x}(t))\leq\frac{\eta}{4}$. Thus, by the Sobolev inequality and (3.24), there exists $N$ such that for any $N\leq i\in\mathbb{N}$,
@@
By (3.22), for any $\eta>0$, there exists $N$ such that for all $i\geq N$, $CI_{t_{0}}(z_{i})\leq\frac{\eta}{2}$. Therefore, one obtains that for $i$ sufficiently large,
@@
and $C_{r}$ is given in Lemma 3.5. As the level sets of $\mathbf{E}^{*}_{L}(\cdot)$ are compact in $E$, we have $K_{\Lambda,-k,0}$ is compact in C([k,0];E)C([-k,0];E) for any $k\in\mathbb{N}$.
@@
We have the following characterisation of $U_{L}$. For simplicity of presentation, we denote
@@
If the infimum is attainable, we write $I_{-\infty}^{\inf}(x)$ as $I_{-\infty}^{\min}(x)$.
@@
where $z$ is the solution of equation (2.3), starting from time $t_{n}$. Since $f\in L^{2}(-\infty,0;H)$, for any $\varepsilon>0$, there exists $T_{\varepsilon}\leq 0$ such that $\frac{1}{2}\|f\|_{L^{2}(-\infty,T_{\varepsilon};H)}^{2}\leq\varepsilon$. Denote $z_{f}^{-t_{n},m_{L}-\psi}$ as the solution of equation (2.3) with $z_{f}^{-t_{n},m_{L}-\psi}(-t_{n})=m_{L}-\psi$. Since $m_{L}-\psi$ is the equilibrium of (2.3) with $f=0$, we have $z_{0}^{-t_{n},m_{L}-\psi}(t)=m_{L}-\psi$ for every $t\in[-t_{n},0]$. Then, by (3.24), for $n\in\mathbb{N}$ sufficiently large there exists $T_{\varepsilon}\in[-t_{n},-t_{n}+1]$ such that
@@
The characterisation of $U_{L}$ allows us to enhance the compactness of the level sets of $I_{-\infty}(\cdot)$ on $C((-\infty,0];E)$ to the compactness of the level sets of $U_{L}(\cdot)$ on $E$. This is the main result of this section and a key result in the theory of large deviations for invariant measures.
@@
such that $U_{L}(x_{n})=I_{-\infty}(z_{n})$. Since $U_{L}(x_{n})\leq r$, we have $\{z_{n}\}\subset K_{-\infty}(r)$. By Proposition 3.6, $K_{-\infty}(r)$ is compact. So, there exists a subsequence $\{z_{n_{k}}\}\subset\{z_{n}\}$ such that $z_{n_{k}}\rightarrow z\in K_{-\infty}(r)$ in $C((-\infty,0];E)$. In particular, $x_{n_{k}}=z_{n_{k}}(0)\rightarrow z(0)$ in $E$. Thanks to Proposition 3.8, we have $U_{L}(z(0))\leq I_{-\infty}(z)\leq r$. That states $z(0)\in\Phi_{L}(r)$. The proof is complete. ∎
4 LDP for invariant measures
In this section, as the main result of this paper, we will prove the LDP for $\{\mu_{\epsilon}\}_{\epsilon>0}$. This is indicated by the LDP of $\{\bar{\mu}_{\epsilon}\}_{\epsilon>0}$, whose relation with $\{\mu_{\epsilon}\}_{\epsilon>0}$ is given by (1.5).